# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 1.3 Overview of this paper
The first term in the loss ${\cal L}^{*}$ are shown to bound $\|f+\Delta v\|_{H^{-1}(\Omega)}$ and the second and third terms are shown to bound $\|g-v\|_{H^{1/2}(\partial\Omega)}$. These in turn give a bound for $\|u-v\|_{H^{1}(\Omega)}$ via (1.3). We go on to show how near minimizers of the suggested loss ${\cal L}^{*}$ give an a priori bound of the approximation error, provided that the approximation method $\Sigma_{n}$ has suitable approximation properties. The a priori guarantee provides a near optimal recovery whenever $\Sigma_{n}$ provides sufficiently good approximation of the elements in the solution model class ${\cal U}$. Notice that this new loss function replaces the first discrete $\ell_{2}$ term in (1.7) by a discrete $\ell_{\gamma}$ loss for a proper choice of $\gamma$ and the second discrete $\ell_{2}$ loss (used to match boundary values) by a weighted $\ell_{2}$ loss.
@@
The remainder of this paper focuses on using minimization over neural network spaces as the numerical method as is the case in PINNs. As noted above, the loss ${\cal L}^{*}$ is the correct loss to use in such a minimization since it is consistent with recovering $u$ in the $H^{1}(\Omega)$. When the loss ${\cal L}^{*}$ is minimized over a neural network space, we call the numerical procedure consistent PINNs (CPINNs).
The later sections of the paper analyze the theoretical gains in using CPINNs over PINNS. We give a priori bounds on how well the solution CPINNs approximates $u$ and, in particular, we provide a sufficient condition on the size $n$ of the neural network to guarantee optimal recovery. These bounds are established under the assumption that the optimization problem is solved (or approximately solved) when optimizing over the NN space $\Sigma_{n}$. As we know, proving that a particular numerical method of optimization (such as gradient decent) converges to a minimizer (or near minimizer) of the loss is a serious issue in optimization, which we do not address here. However, we do prove that for any function $v\in H^{1}(\Omega)$, the quantity ${\cal L}^{*}(v)$ always provides an upper bound of the true error $\|u-v\|_{H^{1}(\Omega)}$ up to the optimal recovery rate for the solution model class ${\cal U}$. This means that ${\cal L}^{*}(v)$ can be used as an a posteriori error estimator for any proposed numerical approximation $v$ to $u$. This error estimate can be used to check whether the output of a PINNs optimization achieves the desired accuracy $\varepsilon$. In other words, although we do not guarantee a priori that PINNs will achieve the accuracy $\varepsilon$ (because of lack of performance analysis of optimization procedures like gradient descent or stochastic gradient descent), we do give an a numerically implementable a posteriori bound on performance which may serve to guarantee optimality.
In summary, a theorem which guarantees that PINNs provide an approximation to $u$ in the $\|\cdot\|_{H^{1}(\Omega)}$ norm with a prescribed accuracy $\varepsilon$ requires not only smoothness assumptions on $f$ and $g$ as described above, assumptions on the spacing and number of elements in the data sites ${\cal X}$ and ${\cal Z}$, and assumptions on the size of $n$, i.e., the complexity of $\Sigma_{n}$, but also an efficient numerical method that properly chooses a good approximation $\hat{u}$ to $u$ in the norm $\|\cdot\|_{H^{1}(\Omega)}$ through minimizing the loss ${\cal L}^{*}$ over $\Sigma_{n}$.
@@ Acknowledgments
We are grateful to Peter Binev, Albert Cohen, Wolfgang Dahmen, and Jinchao Xu for many insightful conversations about the material in this paper.
@@ 2 Besov spaces
We start this section by recalling the definition of Besov spaces and their properties. We confine this paper to the case of functions defined on the domain $\Omega=(0,1)^{d}$, $d\geq 2$. For the range $s>0$, and $0<p,q\leq\infty$, the Besov space $B_{q}^{s}(L_{p}(\Omega))$ is a space of functions with smoothness of order $s>0$ in $L_{p}(\Omega)$. Here $q$ is a secondary index that gives a fine gradation of these spaces. The material in this section is taken for the most part from the papers [18, 16, 17] and the reader will have to refer to those papers for some of the definitions and proofs. Let us also mention that the univariate case is covered in the book [15].
If $r$ is a positive integer and $0<p\leq\infty$ and $f\in L_{p}(\Omega)$, we define the modulus of smoothness $\omega_{r}(f,\cdot)_{p}$ of $f$ by
@@
Let us make some remarks on these spaces. Consider the role of $q$ in this definition. If $q_{2}<q_{1}$, then $B_{q_{2}}^{s}(L_{p}(\Omega))\subset B_{q_{1}}^{s}(L_{p}(\Omega))$, i.e., these spaces get smaller as $q$ gets smaller. Thus, all $B_{q}^{s}(L_{p}(\Omega))$, $q>0$, are contained in $B_{\infty}^{s}(L_{p}(\Omega))$ once $s$ and $p$ are fixed. The effect of $q$ in the definition of the Besov spaces is subtle.
@@ 2.1 The Sobolev spaces $H^{s}(\Omega)$
There are two equivalent ways to describe the space $H^{1/2}(\partial\Omega)$ appearing in A3. The first is as the Besov space $B_{2}^{1/2}(L_{2}(\partial\Omega))$. Here, one needs the concept of Besov spaces on manifolds. In our case, the boundary of $\Omega$ is quite simple since it is the union of the faces of $\Omega$. This definition gives the norm in $H^{1/2}(\partial\Omega)$ through the modulus of smoothness. Using the averaged modulus of smoothness, see the Appendix, one finds that the semi-norm of $g\in H^{1/2}(\partial\Omega)$, is
@@ 2.2 Piecewise polynomial approximation and interpolation
be the tensor product grid of $r^{d}$ equally spaced points in $\overline{\Omega}$ Consider the simplicial (Kuhn-Tucker) decomposition of $\overline{\Omega}$ into simplices $T$,
$\overline{\Omega}=\bigcup\overline{T}.$
Given any one of these simplices $T$, the number $n_{r}$ of grid points in this simplex $\overline{T}$ is the same as the dimension of ${\cal P}_{r}:={\cal P}^{d}_{k}$, where ${\cal P}^{d}_{r}$ denotes the linear space of algebraic polynomials of order $r$ (total degree $r-1$), namely,
@@
Let us consider the standard closed simplex
@@
It is well known that polynomial interpolation using elements of ${\cal P}_{r}$ at the grid points in $G_{r}\cap\overline{T_{0}}$ is well posed (see e.g. Section 4 in [51] or Chapter 2 in [7]. Let $\phi_{j}:=\phi_{j,T_{0}}$, $j=1,\dots,n_{r}$, be the Lagrange polynomial basis for ${\cal P}_{r}$ corresponding to the points in $G_{r}\cap\overline{T}_{0}$. Then, the operator
$L_{T_{0}}(f):=\sum_{{\bf x}_{j}\in G_{r}\cap T_{0}}f({\bf x}_{j})\phi_{j},$ (2.12)
@@
Now, fix $k\geq 0$ and consider the dyadic cubes $I\in{\cal D}_{k}$. If we rescale to any such dyadic cube $I$ and any simplex $T$ in the simplicial decomposition of $I$ and denote by $\phi_{j,T}$ the rescaled polynomial $\phi_{j,T}:=\phi_{j,T_{0}}\circ F_{T}^{-1}$ where $F_{T}:\overline{T}_{0}\rightarrow\overline{T}$ is an affine map from the reference simplex $\overline{T}_{0}$ to $\overline{T}$, we obtain the projector $L_{T}$ onto ${\cal P}_{r}$ which satisfies
@@
Notice that $\Lambda_{r}=\Lambda(T)$. Moreover, the interpolating polynomial $L_{T}(f)$ provides a good approximation to $f$ on $T$. For example, for any polynomial $P\in{\cal P}_{r}$, we have
@@
where ${\cal T}_{k}$ is the collection of all simplicies arising from the (Kuhn-Tucker) decomposition of the dyadic cubes in ${\cal D}_{k}(\overline{\Omega})$ and $\chi_{T}$ is the characteristic function of $T$. The function $S^{*}_{k}(f)$ is continuous on $\overline{\Omega}$ since the number points from $G_{r}$ that are on a shared face of any two simplicies from ${\cal T}_{k}$ is the same as the dimension of the polynomial space ${\cal P}^{d-1}_{r}$
The following theorem (proved in the appendix) describes the appproximation accuracy of $S^{*}_{k}$.
@@
We will also need similar results where the approximation takes place in the $H^{1}(\Omega)=B_{2}^{1}(L_{2}(\Omega))$ norm.
@@
Let $d\geq 2$. If $f\in B_{q}^{s}(L_{p}(\Omega))$ with $s>d/p$ and $0<p\leq 2$ and $S^{*}_{k}(f)$ is as defined in (2.16), then
@@
There are many results concerning the embeddings of Besov spaces into other Besov spaces or $L_{p}$ spaces. We are particularly interested in those when a Besov space is embedded into $C(\Omega)$. For this, we can use the following theorem proved in the appendix
@@
For every $f\in B_{p}^{s}(\Omega)$, $s>d/p$, $0<p\leq\infty$, there is a continuous function $\widetilde{f}$ such that $f=\widetilde{f}$ a.e. In fact, $\widetilde{f}\in$ Lip $\alpha$, with $\alpha:=s-d/p$, i.e.,
@@ 3 Optimal recovery
In this section, we study the question of whether the information $({\bf f},{\bf g})$ at the fixed data sites ${\cal X}:=\{{\bf x}_{1},\ldots,{\bf x}_{\widetilde{m}}\}$ and ${\cal Z}:=\{{\bf z}_{1},\ldots,{\bf z}_{\overline{m}}\}$ is sufficient to determine $u$ to a prescribed accuracy $\varepsilon$ in the $H^{1}(\Omega)$ norm. Questions of this type are well studied and referred to as optimal recovery (OR). The answer to our specific question depends on the numbers $\widetilde{m},\overline{m}$, the positions of the points in ${\cal X}$ and ${\cal Z}$, and the assumptions we make on $f$ and $g$. As we have mentioned, we assume that $f$ and $g$ are continuous so that point evaluation makes sense. It follows that the unknown function $u$ is continuous as well. If we wish to ensure an error bound on how well we can recover $u$ in the $H^{1}(\Omega)$ norm, then the conditions we impose on $f$ and $g$ also have to guarantee that $u$ lies in a compact subset of $H^{1}(\Omega)$. We require that $f,g$ are in certain Besov spaces that compactly embed into the appropriate space of continuous functions, i.e., $f$ lies in a Besov space that compactly embeds into $C(\Omega)$ and $g$ is in a Besov space that compactly embeds into $C(\partial\Omega)$.
Before presenting the results on optimal recovery of this section, let us remark that the results given below are (in essence) all known. For example, the optimal recovery rates for the Besov model classes of this section have all been given in the papers [12, 13, 59] save for one small discrepancy on optimal recovery in $H^{-1}(\Omega)$ (explained below) and for the fact that they do not directly consider recovery on manifolds like $\partial\Omega$. The main distinction between the results in [59] and ours given below is in how the optimal recovery is obtained/proven.The optimal recovery method in [59] uses linear combinations of certain ‘bump functions’, whereas we use continuous piecewise polynomial Lagrange interpolation. One can argue that the fact that optimal recovery can be obtained by interpolation is also known since it is a prominent method of approximation in Finite Element Methods. However, the latter community usually considers Sobolev classes, rather than the more general Besov classes, and also typically restricts $p\geq 1$. In summary, we could not find a direct reference which proves the results given below based on continuous piecewise polynomial interpolation. In any case, those not familiar with optimal recovery may find the results below useful and the OR community can just do a very light reading of this section. The results stated in this section are all proved in detail in the appendix.
We assume that $f,g$ belong to the unit ball of a smoothness space that compactly embeds into the space of continuous functions on the relevant domain. The function $f$ is defined on the domain $\Omega\subset\mathbb{R}^{d}$ with continuous extension on $\overline{\Omega}$. The classical assumptions on function smoothness for a domain $\Omega$ is membership in a Besov space. Accordingly, we shall assume that
@@
where $U({\cal B})$ is the unit ball of ${\cal B}$ and the restriction on $s$ ensures that ${\cal F}$ compactly embeds into the space of continuous functions $C(\overline{\Omega})$.
The function $g$ is defined on the manifold $\partial\Omega$. There are two ways to describe smoothness conditions on such functions $g$. The first is through the trace operator $Tr:=Tr_{\partial\Omega}$. The second is to place smoothness conditions directly on $g$ as a function from $C(\partial\Omega)$. The latter is usually referred to as an intrinsic definition. We choose in this paper to define the set ${\cal G}$ via the trace operator. Namely, we define the model class ${\cal G}$ via the trace of functions from the unit ball of a Besov class $B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega))$ with $\overline{s}>d/\overline{p}$, to ensure continuity. We also place the restrictions $0<\bar{p},\bar{q}\leq 2$ in order to simplify the presentation that follows. In other words, our model class assumptions on the boundary function $g$ take the form
$g\in{\cal G}:=Tr(U(\overline{\cal B})),\ \ \hbox{where}\ \ \overline{\cal B}:=B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega)),\quad\hbox{with}\ \ \bar{s}>d/\bar{p},\quad 0<\bar{p},\bar{q}\leq 2.$ (3.2)
Then ${\cal G}$ is a compact subset of $H^{1/2}(\partial\Omega)$. Note that in general the parameters $s,p,q$ used in defining ${\cal F}$ are different from $\overline{s},\overline{p},\overline{q}$ used in defining ${\cal G}$. These assumptions imply that the function $u$ we want to numerically approximate is an element of the set
@@
Notice that the $g_{i}=u({\bf z}_{i})$, $i=1,\dots,\overline{m}$. Finally, the totality of information we have about the sought after $u$ is that it is in the set
${\cal U}_{\rm data}:={{\cal U}_{{\rm data}(u)}:=}\{\widetilde{u}\in{\cal U}{:\,-\Delta\widetilde{u}({\bf x}_{i})=f_{i},\ i=1,\dots,\widetilde{m},\quad\widetilde{u}({\bf z}_{i})=g_{i},\ \ i=1,\dots,\overline{m}}\}.$
@@
is then the optimal recovery error. We are interested in bounds for $R({\cal U}_{\rm data})_{X}$, where $\|\cdot\|_{X}$ is the norm in which we wish to measure accuracy. We introduce similar notation for the recovery of $f$ and $g$.
@@
where the data values come from any $u\in{\cal U}$. This is measuring the worst possible performance over the class ${\cal U}$ and is called the uniform optimal recovery rate at the fixed data sites $({\cal X},{\cal Z})$. If we prescribe a budget for the number $m=|{\cal X}|+|{\cal Z}|$ of data sites, we can ask for the optimal error we can achieve under such a budget restriction. Accordingly, we define
@@
where the latter is the minimum error we can achieve over all the possible choices of ${\cal X}=\{{\bf x}_{1},\dots,{\bf x}_{\widetilde{m}}\}$. We can similarly ask the question of how well can we can recover $g$ from the data ${\bf g}$, and therefore how well we can recover $u$? Each of these questions has an associated uniform optimal recovery error which is dependent on which norm $\|\cdot\|_{X}$ we use to measure the error of recovery.

### v2

@@ 1.3 Overview of this paper
The first term in the loss ${\cal L}^{*}$ are shown to bound $\|f+\Delta v\|_{H^{-1}(\Omega)}$ and the second and third terms are shown to bound $\|g-v\|_{H^{1/2}(\partial\Omega)}$. These in turn give a bound for $\|u-v\|_{H^{1}(\Omega)}$ via (1.4). We go on to show how near minimizers of the suggested loss ${\cal L}^{*}$ give an a priori bound of the approximation error, provided that the approximation method $\Sigma_{n}$ has suitable approximation properties. The a priori guarantee provides a near optimal recovery whenever $\Sigma_{n}$ provides sufficiently good approximation of the elements in the solution model class ${\cal U}$.
In §8, we analyze a variety of other possible loss functions which provide the same a posteriori error guarantees as the loss ${\cal L}^{*}$. We call the use of any of these loss functions consistent PINNs. If the set $\Sigma_{n}$ is chosen large enough, all these losses will provide a near best optimal recovery (in terms of number of data points) of the solution $u$. We prove that for any function $v\in H^{1}(\Omega)$, the losses suggested in §8 always provide an upper bound of the true error $\|u-v\|_{H^{1}(\Omega)}$ up to the optimal recovery rate for the solution model class ${\cal U}$, provided that the approximant $v$ can be controlled in the $\|\cdot\|_{{\cal U}}$-norm (see Theorem 7.2). This means that these losses can be used as an a posteriori error estimator for any proposed numerical approximation $v$ to $u$. This error estimate can be used to check whether the output of a PINNs optimization achieves the desired accuracy $\varepsilon$.
However, proving that a particular numerical method of optimization (such as gradient descent) converges to a minimizer (or near minimizer) of the loss is a serious issue in optimization, which we do not address here.
@@
Finally, in §10, we numerically test the loss functions described in §8 on a two-dimensional Poisson equation. These results show that using consistent PINNs provides a significant improvement in the solution error, especially for problems with low regularity.
@@ Acknowledgments
We are grateful to Peter Binev, Albert Cohen, Wolfgang Dahmen, and Jinchao Xu for many insightful conversations about the material in this paper. We also thank the referees for their careful reading of the manuscript and their insightful comments/suggestions which helped improve the paper.
@@ 2 Besov spaces
We start this section by recalling the definition of Besov spaces and their properties. For the range $s>0$, and $0<p,q\leq\infty$, the Besov space $B_{q}^{s}(L_{p}(\Omega))$, $\Omega=(0,1)^{d}$, $d\geq 2$, is a space of functions with smoothness of order $s>0$ in $L_{p}(\Omega)$. Here $q$ is a secondary index that gives a fine gradation of these spaces. The material in this section is taken for the most part from the papers [23, 21, 22] and the reader will have to refer to those papers for some of the definitions and proofs. Let us also mention that the univariate case is covered in the book [20].
If $r$ is a positive integer, $0<p\leq\infty$ and $f\in L_{p}(\Omega)$, we define the modulus of smoothness $\omega_{r}(f,\cdot)_{p}$ of $f$ by
@@
Let us make some remarks on these spaces. Consider the role of $q$ in this definition. If $q_{2}<q_{1}$, then $B_{q_{2}}^{s}(L_{p}(\Omega))\subset B_{q_{1}}^{s}(L_{p}(\Omega))$, i.e., these spaces get smaller as $q$ gets smaller. Thus, all $B_{q}^{s}(L_{p}(\Omega))$, $q>0$, are contained in $B_{\infty}^{s}(L_{p}(\Omega))$ once $s$ and $p$ are fixed. The effect of $q$ in the definition of the Besov spaces is subtle. We remark that when $p=q$ and $s$ is not an integer, the space $B^{s}_{p}(L_{p}(\Omega))$ is equivalent to the Sobolev space $W^{s}(L_{p}(\Omega))$ (see [1]).
@@ 2.1 The Sobolev spaces $H^{s}(\Omega)$
There are two equivalent ways to describe the space $H^{1/2}(\partial\Omega)$. The first is as the Besov space $B_{2}^{1/2}(L_{2}(\partial\Omega))$. Here, one needs the concept of Besov spaces on manifolds. In our case, the boundary of $\Omega$ is quite simple since it is the union of the faces of $\Omega$. This definition gives the norm in $H^{1/2}(\partial\Omega)$ through the modulus of smoothness. Using the averaged modulus of smoothness, see (12.3) in the Appendix, §12.1, one finds that the semi-norm of $g\in H^{1/2}(\partial\Omega)$, is
@@ 2.2 Piecewise polynomial approximation and interpolation
be the tensor product grid of $r^{d}$ equally spaced points in $\overline{\Omega}$ Consider the simplicial (Kuhn-Tucker) decomposition of $\overline{\Omega}$ into simplices $T^{0}$,
$\overline{\Omega}=\bigcup\overline{T}^{0}.$ (2.12)
Given any one of these simplices $T^{0}$, the number $n_{r}$ of grid points from $G_{r}$ in the simplex $\overline{T^{0}}$ is the same as the dimension of ${\cal P}_{r}:={\cal P}^{d}_{r}$, where ${\cal P}^{d}_{r}$ denotes the linear space of algebraic polynomials of order $r$ (total degree $r-1$), namely,
@@
Let us consider the closed simplex
@@
It is well known that polynomial interpolation using elements of ${\cal P}_{r}$ at the grid points in $G_{r}\cap\overline{T_{0}}$ is well posed (see e.g. Section 4 in [61] or Chapter 2 in [8]). Let $\phi_{j}:=\phi_{j,T_{0}}$, $j=1,\dots,n_{r}$, be the Lagrange polynomial basis for ${\cal P}_{r}$ corresponding to the points in $G_{r}\cap\overline{T_{0}}$. Then, the operator
$L_{T_{0}}(f):=\sum_{{\bf x}_{j}\in G_{r}\cap\overline{T_{0}}}f({\bf x}_{j})\phi_{j},$ (2.13)
@@
For each $k\geq 0$, we define ${\cal D}_{k}={\cal D}_{k}(\Omega)$ to be the partition of $\Omega$ into dyadic cubes $I$ of side length $2^{-k}$. Here we take a dyadic cube $I\in{\cal D}_{k}$ to be the tensor product of the dyadic intervals $[(j-1)2^{-k},j2^{-k})$, $1\leq j\leq 2^{k}$. We then define the tensor product grid $G_{k,r}$ of $\overline{\Omega}=[0,1]^{d}$,
with spacing $h=2^{-k}(r-1)^{-1}$. We fix $k\geq 0$, $I\in{\cal D}_{k}$, and a simplex $T$ in the simplicial decomposition of $I$, and denote by $\phi_{j,T}$, the rescaled polynomial $\phi_{j,T}:=\phi_{j,T_{0}}\circ F_{T}^{-1}$ where $F_{T}:\overline{T}_{0}\rightarrow\overline{T}$ is an affine map from the reference simplex $\overline{T}_{0}$ to $\overline{T}$. Notice that $\phi_{j,T}$, $j=1,\ldots,n_{r}$, are the Lagrange polynomial basis for ${\cal P}_{r}$, corresponding to the points in $G_{k,r}\cap\overline{T}$, and thus we obtain the projector $L_{T}$ onto ${\cal P}_{r}$ which satisfies
@@
Here $\Lambda_{r}=\Lambda_{r}(T)$, i.e. $\Lambda_{r}(T)$ does not depend on $T$ since for ${\bf x}\in\overline{T}$ we have $\phi_{j,T}({{\bf x}})=\phi_{j,T_{0}}(F^{-1}_{T}({\bf x}))=\phi_{j,T_{0}}({\bf y})$, where ${\bf y}=F^{-1}_{T}({\bf x})\in\overline{T_{0}}$. Moreover, the interpolating polynomial $L_{T}(f)$ provides a good approximation to $f$ on $T$. For example, for any polynomial $P\in{\cal P}_{r}$, we have
@@
where ${\cal T}_{k}$ is the collection of all simplicies arising from the (Kuhn-Tucker) decomposition of the dyadic cubes in ${\cal D}_{k}(\Omega)$ and $\chi_{T}$ is the characteristic function of $\overline{T}$. The function $S^{*}_{k}(f)$ is well defined and continuous on $\overline{\Omega}$ since the number of points from $G_{k,r}$ that are on a shared face of any two simplicies from ${\cal T}_{k}$ is the same as the dimension of the polynomial space ${\cal P}^{d-1}_{r}$
The following theorem (proved in the Appendix, §12.5.1) describes the approximation accuracy of $S^{*}_{k}$.
@@
We will also need similar results where the approximation takes place in the $H^{1}(\Omega)=B_{2}^{1}(L_{2}(\Omega))$ norm. The following theorem, see the Appendix, §12.5.2 for its proof, holds.
@@
Let $d\geq 2$. If $f\in B_{q}^{s}(L_{p}(\Omega))$ with $s>d/p$ and $0<p\leq 2$, and $S^{*}_{k}(f)$ is as defined in (2.18), then
@@
There are many results concerning the embeddings of Besov spaces into other Besov spaces or $L_{p}$ spaces. We are particularly interested in those when a Besov space is embedded into $C(\overline{\Omega})$. For this, we can use the following theorem proved in the Appendix, §12.4.1.
@@
For every $f\in B_{p}^{s}(\Omega)$, $s>d/p$, $0<p\leq\infty$, there is a continuous function $\widetilde{f}$ such that $f=\widetilde{f}$ a.e. In fact, $\widetilde{f}\in$ Lip $\alpha$, with $\alpha:=s-d/p$, that is,
@@ 3 Optimal recovery
In this section, we study the question of whether the information $({\bf f},{\bf g})$ at the fixed data sites
${\cal X}:=\{{\bf x}_{1},\ldots,{\bf x}_{\widetilde{m}}\},\quad{\cal Z}:=\{{\bf z}_{1},\ldots,{\bf z}_{\overline{m}}\},$
is sufficient to determine $u$ to a prescribed accuracy $\varepsilon$ in the $H^{1}(\Omega)$ norm. Questions of this type are well studied and referred to as optimal recovery (OR). The answer to our specific question depends on the numbers $\widetilde{m},\overline{m}$, the positions of the points from ${\cal X}$ and ${\cal Z}$, and the smoothness assumptions we make on $f$ and $g$. Thus, we require that $f,g$ are in certain Besov spaces that compactly embed into $C(\overline{\Omega})$ and $C(\partial\Omega)$, respectively, and are such that $u$ lies in a compact subset of $H^{1}(\Omega)$.
Before presenting the results on optimal recovery in this section, let us remark that the results given below are (in essence) all known. For example, the optimal recovery rates for the Besov model classes of this section have all been given in the papers [13, 14, 69] save for one small discrepancy on optimal recovery in $H^{-1}(\Omega)$ (explained below) and for the fact that they do not directly consider recovery on manifolds like $\partial\Omega$. The main distinction between the results in [69] and ours given below is in how the optimal recovery is obtained/proven. The optimal recovery method in [69] uses linear combinations of certain ‘bump functions’, whereas we use continuous piecewise polynomial Lagrange interpolation. One can argue that the fact that optimal recovery can be obtained by interpolation is also known since it is a prominent method of approximation in Finite Element Methods (see for instance [30]). However, the latter community usually considers Sobolev classes, rather than the more general Besov classes, and also typically restricts $p\geq 1$. In summary, we could not find a direct reference which proves the results given below based on continuous piecewise polynomial interpolation. In any case, those not familiar with optimal recovery may find the exposition below useful and the OR community can just do a very light reading of this section. The statements in this section are all proved in detail in the appendix.
We shall assume that
@@
where $U({\cal B})$ is the unit ball of ${\cal B}$ and the restriction on $s$ ensures that ${\cal F}$ compactly embeds into $C(\overline{\Omega})$.
The function $g$ is defined on the manifold $\partial\Omega$. There are two ways to describe smoothness conditions on such functions $g$. The first is through the trace operator $Tr:=Tr_{\partial\Omega}$. The second is to place smoothness conditions directly on $g$ as a function from $C(\partial\Omega)$. The latter is usually referred to as an intrinsic definition. We choose in this paper to define the set ${\cal G}$ via the trace operator. Namely, we define the model class ${\cal G}$ via the trace of functions from the unit ball of a Besov class $B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega))$ with $\overline{s}>d/\overline{p}$ to ensure continuity. We also place the restrictions $0<\bar{p}\leq 2$ in order to simplify the presentation that follows. In other words, our model class assumptions on the boundary function $g$ take the form
$g\in{\cal G}:=Tr(U(\overline{\cal B})),\ \ \hbox{where}\ \ \overline{\cal B}:=B_{\overline{q}}^{\overline{s}}(L_{\overline{p}}(\Omega)),\quad\hbox{with}\ \ \bar{s}>d/\bar{p},\quad 0<\bar{p}\leq 2.$ (3.2)
Then ${\cal G}$ is a compact subset of both $H^{1/2}(\partial\Omega)$ and $C(\partial\Omega)$. The latter follows because $\overline{\cal B}$ embeds compactly into $C(\overline{\Omega})$ since $\bar{s}>d/\bar{p}$. When taking the trace this compactness is preserved. The former embedding follows since the stated conditions also imply that $\bar{s}>1+(d/\bar{p}-d/2)$ as $d\geq 2$, which implies that $\overline{\cal B}$ embeds into $H^{1}(\Omega)$. Taking the trace will then be compact in $H^{1/2}(\partial\Omega)$. Note that $\bar{s}>1$.
Note that, in general, the parameters $s,p,q$ used in defining ${\cal F}$ are different from $\overline{s},\overline{p},\overline{q}$ used in defining ${\cal G}$. These assumptions imply that the function $u$ we want to numerically approximate is an element of the set
@@
Notice that $g_{i}=u({\bf z}_{i})$, $i=1,\dots,\overline{m}$. Finally, the totality of information we have about the sought after $u$ is that it is in the set
${\cal U}_{\rm data}:={{\cal U}_{{\rm data}(u)}:=}\{\widetilde{u}\in{\cal U}{:\,-\Delta\widetilde{u}({\bf x}_{i})=f_{i},\ i=1,\dots,\widetilde{m},\quad\widetilde{u}({\bf z}_{i})=g_{i},\ \ i=1,\dots,\overline{m}}\},$
@@
is then the optimal recovery error. We are interested in bounds for $R({\cal U}_{{\rm data}(u)})_{X}$, where $\|\cdot\|_{X}$ is the norm in which we wish to measure accuracy. We introduce similar notation for the recovery of $f$ and $g$.
@@
where the data values come from any $u\in{\cal U}$. This is measuring the worst possible performance over the class ${\cal U}$ and is called the uniform optimal recovery rate at the fixed data sites $({\cal X},{\cal Z})$. If we prescribe a budget for the number $m=|{\cal X}|+|{\cal Z}|=\widetilde{m}+\overline{m}$ of data sites, we can ask for the optimal error we can achieve under such a budget restriction. Accordingly, we define
@@
where the latter is the minimum error we can achieve over all the possible choices of data sites ${\cal X}=\{{\bf x}_{1},\dots,{\bf x}_{\widetilde{m}}\}$. We can similarly ask the question of how well can we can recover $g$ from the data ${\bf g}$. Each of these questions has an associated uniform optimal recovery error which is dependent on which norm $\|\cdot\|_{X}$ we use to measure the error of recovery.